# Multi-layer diffusion model of photovoltaic installations

### v1

@@ 2.1 Simulation Details
First, we examine our model with Monte Carlo computer simulations. In order to measure the outcomes at a macroscopic level, we use concentrations, i.e. fractions, of positive adoption states $A$ and positive opinions $S$. See Eq. 4 in Section 3.1 for exact formula. Since $A$ and $S$ are fractions, they belong to the interval $[0,1]$. In all the simulations, we study a system of size $N=2500$ (number of agents, which may be interpreted as a borough in a city). As mentioned in the Introduction, the first layer of the 2-layer network is always a Square Lattice with Moore’s neighborhood [43], and no periodic boundary conditions, which means that almost every agent has exactly 8 neighbors (except those on the edge of the lattice). We denote it as SL($N=2500$, $m=1$), where $m=1$ indicates the range of a neighborhood on a square grid. The second layer is always a 2-dimensional Watts-Strogatz graph [44], with Moore’s neighborhood before rewiring and rewiring probability $\beta=0.2$, and again, without periodic boundary conditions. We wanted a layer that possesses characteristics of a real world social network, but still bears a resemblance to the other layer. We denote it as WS2D($N=2500$, $m=1$, $\beta=0.2$). We study the behavior of the system with respect to the following parameters:
  • 1.
    probability of independence $p$ (probability that an agent changes opinion independently of social influence),
  • 2.
    probability of adoption $a_{1}$ (probability that an agent with a positive opinion changes its adoption state to a positive),
  • 3.
    probability of loosing adoption $a_{2}=ha_{1}$ (probability that an agent with a negative opinion changes its adoption state to a negative).
Since all these are probabilities, they all belong to the interval $[0,1]$, as does the $h$ coefficient in our study. Although the size of the group of influence $q$ is a parameter as well, here we keep it constant, $q=4$, as its impact in the $q$-voter model with independence is well studied [39]. This will help us to reduce computational efforts slightly. In all the simulations, we start from a fully unadopted, negative state, i.e., $A(0)=0$, $S(0)=0$. Lastly, for illustrative purposes, the time horizon $T$ for each simulation is limited to $5000$ Monte Carlo steps (MCS), and the number of time trajectories to $10$.
@@ 2.2 Simulation Model
Figure 2: 10 simulated time trajectories of $A$ and $S$ for different values of $p$: AND variant (top), OR variant (middle). Values of $a_{1}$: $a_{1}=0.04$ (left), $a_{1}=0.16$ (right). First layer - SL(N,1), second layer - WS2D(N,1,0.2), size $N=2500$ and $a_{2}=0.5a_{1}$ in all cases. Bottom row is a just slice for a single value of $p=0.1$ from the top for illustrative purposes.
Let us begin with Fig. 2. It shows evolution of the system, for both variants (AND and OR), and a range of values of $p$, $a_{1}$ and $a_{2}$, where $a_{2}=0.5a_{1}$ in all cases. We can observe that several final outcomes (so-called exit states) are possible. When $p$ is low, we retain an unadopted state ($A(T)\approx 0$, $S(T)\approx 0$). None of the agents have a positive opinion or adoption state, save a few rebels. Then, there is a critical value of $p$, above which the system becomes adopted ($A(T)\approx 1$, $S(T)\approx 1$). If the value of $p$ is high, independence surpasses conformity and we end up with a disordered state ($A(T)\approx 0.5$, $S(T)\approx 0.5$). There is a visible difference between the two variants: AND and OR. In the first, much lower values of $p$ are required to enter adopted or disordered state than in the latter. Finally, the value of $a_{1}$ itself has no impact on the exit state (except for $a_{1}=0$), only on the time needed to reach it (exit time), with the higher value speeding up the process. If $a_{1}=0$, the system cannot leave the unadopted state.
@@
Figure 3: 10 simulated time trajectories of $A$ and $S$ for different values of $p$: AND variant (top) and OR variant (middle). Values of $a_{1}$: $a_{1}=0.04$ (left), $a_{1}=0.16$ (right). First layer - SL(N,1), second layer - WS2D(N,1,0.2), size $N=2500$ and $a_{2}=0.25a_{1}$ in all cases.
What affects the exit state, however, is the relationship between $a_{1}$ and $a_{2}$, namely the coefficient $h$. Let us move to Fig. 3. There, $a_{2}$ is twice as small relative to $a_{1}$ as before, $a_{2}=0.25a_{1}$. With this decrease in $a_{2}$ much lower values of $p$ are needed, in both variants, for the system to become adopted. This itself is a rather trivial conclusion, as now the probability of loosing adoption is four times smaller that the one of adopting. There is, however, a secondary effect to it, due to the fact how the first layer of the network impacts opinions in our model. Positive adoption states support positive opinions (and vice versa). For this reason, the adopted state is not only easier to reach, but also more difficult to disorder.
@@
Figure 4: 10 simulated time trajectories of $A$ and $S$ for different values of $p$: AND variant (top) and OR variant (middle). Values of $a_{1}$: $a_{1}=0.04$ (left), $a_{1}=0.16$ (right). First layer - SL(N,1), second layer - WS2D(N,1,0.2), size $N=2500$ and $a_{2}=0$ in all cases.
If we move forward to Fig. 4, we observe this phenomenon to be even more enhanced. With the probability of loosing adoption $a_{2}=0$, the system always escapes from the unadopted state, except when $a_{1}=0$ as well. Because adoption can never be lost, only opinion can disorder with an increase of independence $p$.
@@
We decided not to show a symmetrical case $a_{1}=a_{2}$, because there is no adopted state there. The system can only evolve into total disorder or remain unadopted.
@@ 2.3 Mean-Field Approximation
Monte Carlo computer simulations are the accurate method when it comes to agent-based models. Unfortunately, as mentioned in the Introduction, they can be very time-consuming. For this reason, we derive a Mean-Field Approximation for our model. The equations describing the dynamical system are derived in Section 3.2 (Eqs. 5-13). Here, we recall them:
$\frac{\partial A}{\partial t}$ $=S(1-A)a_{1}-(1-S)Aa_{2},$ (1)
$\frac{\partial S}{\partial t}$ =$\stackrel{{\scriptstyle\textbf{AND}}}{{=}}(1-S)\left(\frac{1}{2}p+(1-p)S^{q}A^{q}\right)-S\left(\frac{1}{2}p+(1-p)(1-S)^{q}(1-A)^{q}\right),$ (2)
$\frac{\partial S}{\partial t}$ =$\stackrel{{\scriptstyle\textbf{OR}}}{{=}}(1-S)\left(\frac{1}{2}p+(1-p)\left(S^{q}\sum_{i}^{q-1}\binom{q}{i}A^{i}(1-A)^{q-i}\right.\right.$
$\left.\left.+A^{q}\sum_{i}^{q-1}\binom{q}{i}S^{i}(1-S)^{q-i}+S^{q}A^{q}\right)\right)$
$-S\left(\frac{1}{2}p+(1-p)\left((1-S)^{q}\sum_{i}^{q-1}\binom{q}{i}(1-A)^{i}A^{q-i}\right.\right.$
$\left.\left.+(1-A)^{q}\sum_{i}^{q-1}\binom{q}{i}(1-S)^{i}S^{q-i}+(1-S)^{q}(1-A)^{q}\right)\right).$ (3)
The first equation, $\frac{\partial A}{\partial t}$, remains the same both cases. Only the formulas for $\frac{\partial S}{\partial t}$ differ between AND and OR variants. Thanks to MFA, we are can examine a wide range of parameter values, which we could not achieve with Monte Carlo simulations in any reasonable amount of time.
@@
Figure 5: 10 simulated time trajectories of $A$ and $S$ versus numerically obtained time trajectory, for different values of $p$ and $a_{1}$. AND variant (top 4) and OR variant (bottom 4). First layer - SL(N,1), second layer - WS2D(N,1,0.2), size $N=2500$ and $a_{2}=0.5$ in all cases.
@@
Before we move on to the MFA results themselves, we should check if the approximation is any good. For this, see Fig. 5, where we compare simulation results against MFA. As shown there, mean-field gives fairly good approximation, except for the bottom row (OR variant, $p=0.2$). This is due to the fact, that in MFA critical values of $p$ required for the system to adopt are slightly higher, as presented in the following figures.
@@
Figure 6: Mean-Field exit states (left) and exit times (right) for AND variant (top) and OR (bottom). Left side compares MFA numerical results (markers) vs analytical (continuous lines). Numerical results cover only a portion of analytical ones, as they present exit states from a single pair of initial conditions $(A_{0}=0,S_{0}=0)$ only, while the latter show all the possible exit states. Right side shows exit times obtained with numerical methods. Adoption probabilities $a_{1}=0.5$ (left side only) and $a_{2}=0.5a_{1}$.
@@
Mean-field time trajectories, and hence the exit times, are obtained numerically from Eqs. 1-3. Analytically, we are only able to compute the exit states (see Section 3.2, Eqs. 14-19). We compare the two in Fig. 6, for $a_{2}=0.5a_{1}$. Numerically computed exit states perfectly match analytical ones. Segments not covered by the numerical results are due to the fact that the analytical solution shows all the possible exit states, while the numerical one only those achievable from given initial conditions ($A(0)=0$, $S(0)=0$). Clearly visible here are the three possible exit states: unadopted, adopted and disordered, although a transition between the latter two in OR variant is very smooth. As already mentioned, the value of $a_{1}$ itself has no impact on the exit state (except for $a_{1}=0$). Therefore, we set it arbitrarily to $0.5$ (only left side of the figure). On the right, exit times with respect to $a_{1}$ and $p$ are presented. Exit times drop dramatically with an increase of $a_{1}$, but only for very low values of $a_{1}$. After that, the change is unnoticeable. There are significant ,,ridges” for the values of $p$ corresponding to transitions between exit states (left size). There are two such ridges in AND variant, but only one in OR variant, as transition between adopted and disordered is very smooth there. These increases in exit times are logical, the system needs more time to ,,decide” which path to take, and consistent with our knowledge on phase transitions.
Figure 7: Mean-Field exit states (left) and exit times (right) for AND variant (top) and OR (bottom). Left side compares MFA numerical results (markers) vs analytical (continuous lines). Adoption probabilities $a_{1}=0.5$ (left side only) and $a_{2}=0.25a_{1}$. The rest is analogous to Fig. 6.
Figure 8: Mean-Field exit states (left) and exit times (right) for AND variant (top) and OR (bottom). Left side compares MFA numerical results (markers) vs analytical (continuous lines). Adoption probabilities $a_{1}=0.5$ (left side only) and $a_{2}=0$. The rest is analogous to Fig. 6.
@@
When we decrease $a_{2}$ to $a_{2}=0.25a_{1}$ (see Fig. 7), transitions between adopted and disordered exit states become less apparent in both variant of the model. Moreover, values of $p$ needed for adoption decrease, while those needed for disorder increase, which is consistent with the simulation results. Finally, we study the $a_{2}=0$ case with MFA (Fig. 8). Again, the results are consistent with the simulations. We observe that only for $a_{1}=0$ unadopted state can be retained. For any $a_{1}>0$ we always end up with total adoption ($A=1$), while the opinion $S$ slowly transitions to the disorder as $p$ increases.
3 Model and Methods
Figure 9: Graphical representation of the model. The network consists of 2 layers: 1 – Square Lattice (SL), on which adoption states $A_{i}$ are visible, and 2 – 2-dimensional Watts-Strogatz (WS2D) with opinions $S_{i}$. Adoption states $A_{i}$ are represented by outer circles (green – $A_{i}=+1$, red – $A_{i}=-1$), while opinions – by inner circles. Grey areas correspond to adoption states or opinions unknown to the target agent (marked with a dark blue circle). Groups of influence (of size $q=4$, marked with light blue circles) are constructed independently on each layer. In the given example, such a choice would be sufficient to change target’s opinion ($S_{target}=-1\rightarrow+1$) in OR variant, but not in AND variant, as unanimity is only achieved in one of the two groups of influence.
3.1 Simulation Model
We consider a set of $N$ agents, each of which is characterized by 2 binary variables: adoption $A_{i}=\pm 1$ (adopted or not adopted), and opinion $S_{i}=\pm 1$ (positive or negative), for $i=1,2,\dots,N$. Agents are located on a 2-layer network. The first layer of the network depicts the spacial location of an agent and is represented by a square lattice (SL) with Moore’s neighborhood [43]. At this level, an agent sees only the adoption states $A$ of its neighbors. The second layer corresponds to structure of social ties of agents, which is described by a 2-dimensional Watts-Strogatz graph [44]. Here, an agent sees only the opinions $S$ of its neighbors.
We investigate the model through Monte Carlo simulations with a random sequential updating scheme. Within a single simulation, the time is measured in so-called Monte Carlo steps (MCS). One MCS consists of $N$ elementary events, each divided into the following substeps of length $\Delta t=\frac{1}{N}$:
  • 1.
    Select a target agent $i$ randomly (uniformly from $N$ nodes).
  • 2.
[…]
  • 3.
    With the complementary probability $1-p$ (if $r_{1}>p$), 2 groups of influence are constructed (one from each layer):
    • (a)
      Randomly select $q$ neighbors of agent $i$ from the first layer, $j_{11}$, $j_{12}$, $\dots$, $j_{1q}$, with possibility of repetition.
    • (b)
      Randomly select $q$ neighbors of agent $i$ from the second layer, $j_{21}$, $j_{22}$, $\dots$, $j_{2q}$, with possibility of repetition.
  • 4.
    AND variant If both groups of influence are unanimous within and between each other, $\sum_{i=1}^{q}A_{j_{1i}}(t)=\sum_{i=1}^{q}S_{j_{2i}}(t)=\pm q$, the agent $i$ conforms and changes its opinion, i.e., $S_{i}(t+\Delta t)\leftarrow A_{j_{11}}(t)$. Otherwise, nothing happens.
  • 5.
    OR variant If at least one group of influence is unanimous and the other is not unanimous oppositely, $\sum_{i=1}^{q}A_{j_{1i}}(t)=\pm q\neq-\sum_{i=1}^{q}S_{j_{2i}}(t)$, the agent $i$ conforms to the unanimous group, $S_{i}(t+\Delta t)\leftarrow A_{j_{11}}(t)$, or $S_{i}(t+\Delta t)\leftarrow S_{j_{21}}(t)$ when $\sum_{i=1}^{q}S_{j_{2i}}(t)=\pm q\neq-\sum_{i=1}^{q}A_{j_{1i}}(t)$. Otherwise, nothing happens.
  • 6.
    Finally, the agent $i$ adjusts its adoption state $A_{i}$ to its opinion $S_{i}$ with some probability. Draw a random number $r_{2}\sim U(0,1)$. Then:
    • (a)
      If the agent has positive opinion $S_{i}(t)=1$ but a negative adoption state $A_{i}(t)=-1$, it adopts with probability $a_{1}$ (if $r_{2}<a_{1}$), $A_{i}(t+\Delta t)\xleftarrow[]{a_{1}}1$.
    • (b)
      Analogously, if the agent has negative opinion $S_{i}(t)=-1$ but a positive adoption state $A_{i}(t)=1$, it looses its adoption with probability $a_{2}=ha_{1}$ (if $r_{2}<a_{2}$), $A_{i}(t+\Delta t)\xleftarrow[]{a_{2}}-1$.
See Fig. 9 for a graphical representation of the model.
To clarify, we do the following. At the beginning of each independent simulation (trajectory), we set the initial conditions. In simulations, we always start from a fully unadopted, negative state, i.e., $\forall_{i}A_{i}(0)=-1,S_{i}(0)=-1$. Then, we perform Monte Carlo steps until a determined time horizon $T$ is reached ($T$ is the number of MCS). We repeat independent simulations multiple times with the same set of parameter values for a better statistical accuracy. We also run separate simulations for different sets of parameter values.
To examine the model on a macroscopic scale, we use two measures, concentration, i.e., fraction, of positive adoption states and concentration positive opinions:
For the sake of simplicity, we refer to them as just $A(A=c_{A})$ and $S(S=c_{S})$.
3.2 Mean-Field Approximation
Here, we derive a set of equations describing the dynamical system in Mean-Field Approximation (Eqs. 1-3). First, let us assume that the system is homogeneous, i.e., every agent possesses the same knowledge of the entire system. Then, we can write down probabilities that the concentrations of positive adoption states/opinions will increase or decrease, the so-called transition probabilities. These are:
$\gamma_{A,AND}^{+}$ $=S(1-A)a_{1},$ (5)
$\gamma_{A,AND}^{-}$ $=(1-S)Aha_{1},$ (6)
$\gamma_{S,AND}^{+}$ $=(1-S)\left(\frac{1}{2}p+(1-p)S^{q}A^{q}\right),$ (7)
$\gamma_{S,AND}^{-}$ $=S\left(\frac{1}{2}p+(1-p)(1-S)^{q}(1-A)^{q}\right),$ (8)
in AND variant, and
$\gamma_{S,OR}^{+}$ $=(1-S)\left(\frac{1}{2}p+(1-p)\left(S^{q}\sum_{i}^{q-1}\binom{q}{i}A^{i}(1-A)^{q-i}\right.\right.$
$\left.\left.+A^{q}\sum_{i}^{q-1}\binom{q}{i}S^{i}(1-S)^{q-i}+S^{q}A^{q}\right)\right),$ (9)
$\gamma_{S,OR}^{-}$ $=S\left(\frac{1}{2}p+(1-p)\left((1-S)^{q}\sum_{i}^{q-1}\binom{q}{i}(1-A)^{i}A^{q-i}\right.\right.$
$\left.\left.+(1-A)^{q}\sum_{i}^{q-1}\binom{q}{i}(1-S)^{i}S^{q-i}+(1-S)^{q}(1-A)^{q}\right)\right),$ (10)
in OR variant. Probabilities $\gamma_{A}^{+}$ and $\gamma_{A}^{-}$ remain the same in both variants. i.e., $\gamma_{A,AND}^{+}=\gamma_{A,OR}^{+}$ and $\gamma_{A,AND}^{-}=\gamma_{A,OR}^{-}$. Then, the dynamical system is described by a pair of differential equations:
$\frac{\partial A}{\partial t}$ $=\gamma_{A,AND}^{+}-\gamma_{A,AND}^{-}=\gamma_{A,OR}^{+}-\gamma_{A,OR}^{-},$ (11)
[…]
In order to find exit states, i.e. values of $A$ and $B$ for which the system can no longer change, we equate the above to zero:
$\frac{\partial A}{\partial t}=0,\,\,\,\,\frac{\partial S}{\partial t}=0.$ (14)
First, we express $A$ as a function of $S$:
$A=\frac{S}{S+h-Sh}\,\,\,\,\text{if}\,\,\,\,a_{1}\neq 0.$ (15)
As shown, $a_{1}$ has no impact on exit states when $a_{1}\neq 0$. If $a_{1}=0$, then $\forall_{A}\frac{\partial A}{\partial t}=0$. Now that we expressed $A$ with $S$, we must find relationship between $S$ and $p$. Unfortunately, finding $S$ as a function of $p$ is impossible due to the arbitrary value of $q$. Hence, we derive the opposite relation that satisfies Eq. 14. It is:
$p=\frac{S^{q}A^{q}-SS^{q}A^{q}-S(1-S)^{q}(1-A)^{q}}{S^{q}A^{q}-SS^{q}A^{q}-S(1-S)^{q}(1-A)^{q}+S-\frac{1}{2}},$ (16)
in AND variant, and
$p=\frac{\sum_{A}^{+}\sum_{S}^{+}-S\sum_{A}^{+}\sum_{S}^{+}-S\sum_{A}^{-}\sum_{S}^{-}}{\sum_{A}^{+}\sum_{S}^{+}-S\sum_{A}^{+}\sum_{S}^{+}-S\sum_{A}^{-}\sum_{S}^{-}+S-\frac{1}{2}},$ (17)
in OR variant, where
$\sum_{A}^{+}\sum_{S}^{+}$ $=\left(S^{q}\sum_{i}^{q-1}\binom{q}{i}A^{i}(1-A)^{q-i}+A^{q}\sum_{i}^{q-1}\binom{q}{i}S^{i}(1-S)^{q-i}+S^{q}A^{q}\right),$ (18)
$\sum_{A}^{-}\sum_{S}^{-}$ $=\left((1-S)^{q}\sum_{i}^{q-1}\binom{q}{i}(1-A)^{i}A^{q-i}\right.$
$\left.+(1-A)^{q}\sum_{i}^{q-1}\binom{q}{i}(1-S)^{i}S^{q-i}+(1-S)^{q}(1-A)^{q}\right).$ (19)
Finally, with $p$ and $A$ dependent on $S$, we can draw exit states of $A$ and $S$ as a function of $p$, what we did in Section 2.3.

### v2

@@ 3.1 Simulation Details
First, we examine our model with Monte Carlo computer simulations. In order to measure the outcomes at a macroscopic level, we use concentrations, i.e. fractions, of positive adoption states $c_{A}$ and positive opinions $c_{S}$, see Definition 2. In all the simulations, we study a system of size $N=2500$ (number of agents, which may be interpreted as a borough in a city). As mentioned in the Introduction, the first layer of the two-layer network is always a Square Lattice with Moore’s neighborhood [44], and no periodic boundary conditions, which means that almost every agent has exactly 8 neighbors (except those on the edges of the lattice). We denote it as SL($N=2500$, $m=1$), where $m=1$ indicates the range of a neighborhood on a square grid. The second layer is always a two-dimensional Watts-Strogatz graph [45], with Moore’s neighborhood before rewiring and rewiring probability $\beta=0.2$, and again, without periodic boundary conditions. We wanted a layer that possesses characteristics of a real world social network, but still bears a resemblance to the other layer. We denote it as WS2D($N=2500$, $m=1$, $\beta=0.2$). We study the behavior of the system with respect to the parameters $p$, $a_{1}$ and $a_{2}$. Although the size of the group of influence $q$ is a parameter as well, here we keep it constant, as its impact in the $q$-voter model with independence is well studied [40]. We choose $q=4$ motivated by real life experiments [50]. This will help us to reduce computational efforts slightly. In all the simulations, we start from a fully unadopted, negative state, i.e., $c_{A}(0)=0$, $c_{S}(0)=0$. Lastly, for illustrative purposes, the time horizon $T$ for each simulation is limited to $5000$ Monte Carlo steps (MCS), and the number of time trajectories to $10$.
@@ 3.2 Simulation Model
Figure 3: 10 simulated time trajectories of $c_{A}$ (red) and $c_{S}$ (blue) for different values of $p$: the AND variant (top), the OR variant (bottom). Values of $a_{1}$: $a_{1}=0.04$ (left), $a_{1}=0.16$ (right). First layer – SL(N,1), second layer – WS2D(N,1,0.2), size $N=2500$ and $a_{2}=0.5a_{1}$ in all cases.
Let us begin with Fig. 3. It shows evolution of the system, for both variants (AND and OR), and a range of values of $p$, $a_{1}$ and $a_{2}$, where $a_{2}=0.5a_{1}$ in all cases. We can observe that several final outcomes are possible. When $p$ is low, we retain an unadopted state ($c_{A}(T)\approx 0$, $c_{S}(T)\approx 0$). None of the agents have a positive opinion or adoption state, except for a few rebels. Then, there is a critical value of $p$, above which the system becomes adopted ($c_{A}(T)\approx 1$, $c_{S}(T)\approx 1$). If the value of $p$ is high, independence surpasses conformity and we end up with a disordered system ($c_{A}(T)\approx 0.5$, $c_{S}(T)\approx 0.5$). There is a visible difference between the two variants: AND and OR. In the first, much lower values of $p$ are required to enter adopted or disordered state than in the latter. Finally, the value of $a_{1}$ itself has no impact on the final state (except for $a_{1}=0$, which is not shown here), only on the time needed to reach it, with the higher value speeding up the process.
@@
Figure 4: 10 simulated time trajectories of $c_{A}$ (red) and $c_{S}$ (blue) for different values of $p$: the AND variant (top), the OR variant (bottom). Values of $a_{1}$: $a_{1}=0.04$ (left), $a_{1}=0.16$ (right). First layer – SL(N,1), second layer – WS2D(N,1,0.2), size $N=2500$ and $a_{2}=0.25a_{1}$ in all cases.
What affects the final state, however, is the relationship between $a_{1}$ and $a_{2}$, namely the coefficient $h$ (Eq. (1)). Let us move to Fig. 4. There, $a_{2}$ is twice as small relative to $a_{1}$ as before, $a_{2}=0.25a_{1}$. With this decrease in $a_{2}$ much lower values of $p$ are needed, in both variants, for the system to become adopted. This itself is a rather trivial conclusion, as now the probability of loosing adoption is four times smaller that the one of adopting. There is, however, a secondary effect to it, due to the fact how the first layer of the network impacts opinions in our model. Positive adoption states support positive opinions (and vice versa). For this reason, the adopted state is not only easier to reach, but also more difficult to disorder.
@@
We decided not to examine a symmetrical case $a_{1}=a_{2}$, i.e. $h=1$, because there is no adopted state there. From $c_{A}(0)=0$, $c_{S}(0)=0$, the system can only evolve into total disorder or remain unadopted.
@@ 3.3 Mean-Field Approximation
Monte Carlo computer simulations are the first choice when it comes to agent-based models. Unfortunately, as mentioned in the Introduction, they can be very time-consuming. For this reason, we derive a set of equations describing dynamics of the system (Eqs. (3)-(5)), under the assumption that each layer of the network is a complete graph of size $N\to\infty$. Thanks to this approach, we can examine a wide range of parameter values, which we could not achieve with Monte Carlo simulations in any reasonable amount of time.
@@
Figure 5: 10 simulated time trajectories of $c_{A}$ and $c_{S}$ versus numerically obtained time trajectory (from Eqs. (3)-(5)), for different values of $p$ and $a_{1}$. the AND variant (top 4) and the OR variant (bottom 4). First layer – SL(N,1), second layer – WS2D(N,1,0.2), size $N=2500$ and $a_{2}=0.5a_{1}$ in all cases.
@@
Before discussing the analytical results themselves, we compare them with the simulations. For this, see Fig. 5, where we combine the latter and numerically obtained time trajectories from Eqs. (3)-(5). As shown there, mean-field gives fairly good approximation, except for the bottom row (OR variant, $p=0.2$). This is due to the fact, that in the MFA critical values of $p$ required for the system to adopt are slightly higher, as presented in the next two Figures (Figs. 6 and 7).
@@
Figure 6: Stationary states (left) and time to reach a stationary state (right) for the AND variant (top) and the OR (bottom). Left side compares numerical results (from Eqs. (3)-(5); markers) vs analytical (from Eqs. (16)-(18); continuous lines). Numerical results cover only a portion of analytical ones, as they present stationary states from a single pair of initial conditions $(c_{A}(0)=0,c_{S}(0)=0)$ only, while the latter show all the possible stationary states. Right side shows time to reach a stationary state obtained with numerical methods. Adoption probabilities $a_{1}=0.5$ (left side only) and $a_{2}=0.5a_{1}$.
@@
Mean-field time trajectories, and hence times to reach a stationary state, are obtained numerically from Eqs. (3)-(5). Analytically, we are only able to compute stationary states (Eqs. (16)-(18)). We compare the two in Fig. 6, for $a_{2}=0.5a_{1}$. Numerically computed stationary states perfectly match analytical ones. Segments not covered by the numerical results are due to the fact that the analytical solution shows all the possible stationary states, while the numerical one only those achievable from given initial conditions ($c_{A}(0)=0$, $c_{S}(0)=0$). Clearly visible here are the three possible groups of stationary states: unadopted ($c_{A}(T)\approx 0$, $c_{S}(T)\approx 0$), adopted ($c_{A}(T)\approx 1$, $c_{S}(T)\approx 1$) and disordered ($c_{A}(T)\approx 0.5$, $c_{S}(T)\approx 0.5$), although a transition between the latter two in the OR variant is very smooth. As already mentioned, the value of $a_{1}$ ($a_{1}\in(0,1]$) itself has no impact on a stationary state. Therefore, we set it arbitrarily to $0.5$ (only left side of Fig. 6). On the right, times to reach a stationary state with respect to $a_{1}$ and $p$ are presented. These drop dramatically with an increase of $a_{1}$, but only for very low values of $a_{1}$. After that, the change is unnoticeable. There are significant ,,ridges” for the values of $p$ corresponding to transitions between groups of stationary states (left side). There are two such ridges in the AND variant, but only one in the OR variant, as transition between adopted and disordered is very smooth there. These increases in times are logical, as the system needs more time to ,,decide” which path to take, and consistent with our knowledge on phase transitions.
Figure 7: Stationary states (left) and time to reach a stationary state (right) for the AND variant (top) and the OR (bottom). Left side compares numerical results (from Eqs. (3)-(5); markers) vs analytical (from Eqs. (16)-(18); continuous lines). Numerical results cover only a portion of analytical ones, as they present stationary states from a single pair of initial conditions $(c_{A}(0)=0,c_{S}(0)=0)$ only, while the latter show all the possible stationary states. Right side shows time to reach a stationary state obtained with numerical methods. Adoption probabilities $a_{1}=0.5$ (left side only) and $a_{2}=0.25a_{1}$.
@@
When we decrease $a_{2}$ to $0.25a_{1}$ (see Fig. 7), transitions between adopted and disordered states become less apparent in both variants of the model. Moreover, values of $p$ needed for adoption decrease, while those needed for disorder increase, which is consistent with the simulations results.